40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Sep 20 11:39:04 2019, gautam, Update, Optical Levers, ETMX Oplev HeNe Dead ETMX_OLdead.png
    Reply  Sun Sep 29 16:08:25 2019, gautam, Update, Optical Levers, ETMX Oplev HeNe replaced OLRIN_20190929.pdf
Message ID: 14911     Entry time: Sun Sep 29 16:08:25 2019     In reply to: 14902
Author: gautam 
Type: Update 
Category: Optical Levers 
Subject: ETMX Oplev HeNe replaced 

To facilitate POX locking investigations, I replaced this HeNe today with one of the spares Chub/Steve had acquired some time ago. Details:

  • Part number: Lumentum 22037130 (1103P)
  • Serial number: PA00836
  • Manufacture date: 01/2019
  • Power output: ~2.64 mW (Measured with Ophir power meter in the 632nm setting)
  • Power received on QPD: ~0.37 mW = ~18700 cts (Measured with Ophir power meter in the 632nm setting)

The RIN of the sum channel with the Oplev servo engaged, along with that for the other core FPMI optics, in shown in Attachment #1. The ETMX HeNe RIN is compatible with the other HeNes in the lab (the high-frequency behaviour of the BS Oplev is different from the other four because the QPD whitening electronics are different).

Not sure what to make of the ETMY RIN profile being so different from the others, seems like some kind of glitchy behaviour, I could see the mean level of the ASD moving up and down as I was taking the averages in DTT. Needs further investigation.

The old / broken HeNe is placed i(nside the packaging of the abovementioned replacement HeNe) on Steve's old desk for disposal in the proper way.

*It looked like Steve had hooked up a thermocouple to be able to monitor the temperature of the HeNe head. I removed this feature as I figured if we don't have this hooked up to the DAQ, it isn't a really useful diagnostic. If we want, we can restore this in a more useful way.

Quote:

While working on recovering interferometer alignment, I noticed that the ETMX Oplev SUM channel reported 0 counts. Attachment #1 shows the 200 day trend - despite the missing data, the accelerating downward decay is evident. I confirmed that there is no light coming out of the HeNe by walking down to EX. The label on the HeNe says it was installed in March 2017, so the lifetime was ~30 months. Seems a little short? I may replace this later today.

Attachment 1: OLRIN_20190929.pdf  69 kB  Uploaded Sun Sep 29 17:25:43 2019  | Hide | Hide all
OLRIN_20190929.pdf
ELOG V3.1.3-